Need help in subracting and adding decimals
1) 3.18 - 5.72 - 6.4 = ?
2) -8.9 + 7.36 - 14.2 = ?
3) What is 4.793 less than -6.82 ?
4) -125.41 + 361.55 =?
I have answers but I just want to check with your answers.
Thank you

Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

1) 3.18 - 5.72 - 6.4 = ?
= 3.18 - 12.12 (adding 5.72 and 6.4)
= - 8.94
--------------
2) -8.9 + 7.36 - 14.2 = ?
= 7.36 - 23.1 (adding the negative terms)
= 15.74
3) What is 4.793 less than -6.82 ?
That makes no sense. Should it say "less", rather than "less than"?
---------------
4) -125.41 + 361.55 =?
Why don't you do that one.
-----------------------

Answer by jim_thompson5910(35256)

1) 3.18 - 5.72 - 6.4 = -2.54 - 6.4 = -8.94


So the answer is -8.94



2) -8.9 + 7.36 - 14.2 = -1.54 - 14.2 = -15.74


So the answer is -15.74


3)


"4.793 less than -6.82" translates to -6.82-4.793 and simplifies to

-6.82-4.793=-11.613


So the answer is -11.613


4)


-125.41 + 361.55 = 236.14


So the answer is 236.14